OpenStudy (awesome781):

A bullet with a mass of 5.00 x 10^-3 kg is loaded into a gun. The loaded gun has a mass of 0.52 kg. The bullet is fired, causing the empty gun to recoil at a speed of 2.1 m/s. What is the speed of the bullet?

OpenStudy (anonymous):

You can use conservation of momentum \[p_i = p_f\] noting that for each object in the problem \[p=mv\] Initial momentum is zero, and there are two terms for final momentum of the system - one for the bullet, and one for the empty gun (note that the problem only gives you the mass of the *loaded* gun, so you have to find the mass of the gun without a bullet in it). The unknown that you solve for is the velocity of the bullet.

OpenStudy (anonymous):

220 m/s mass of the bullet * velocity of the bullet = mass of the gun * velocity of the gun 0.005 kg * v = 0.52 kg * 2.1 m/s -> v = (0.52kg * 2.1 m/s) / 0.005 kg = 218.4 m/s

OpenStudy (anonymous):

Remember both that we're looking for the velocity of the bullet, and that the mass of the gun is actually \[m_{gun} = .515kg\] when the bullet has left the chamber. The gun did indeed recoil with a velocity of 2.1m/s, so the velocity of the bullet should be much higher. \[v_{bullet} = \frac{m_{gun} v_{gun}}{m_{bullet}} = \frac{(.515kg)(2.1m/s)}{.005kg} = 216 m/s\]

OpenStudy (loser66):

oh yea, @AllTehMaffs is right, hehehe. My bad, I am terribly sorry @lillybye

OpenStudy (loser66):

let me take off my stupid comments, hehehehe . Again, I apology @lillybae